40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 18:03:47 2014, ericq, Update, Green Locking, uPDH Box Checkup uPDH_X.JPGuPDH_Y.JPGupdhTFs.pdfupdhNoises.pdfuPDH_Aug27.zip
    Reply  Thu Aug 28 17:34:20 2014, ericq, Update, Green Locking, uPDH Box Checkup TEK00006.PNG
       Reply  Fri Aug 29 17:28:07 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Fri Aug 29 23:53:35 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Tue Sep 2 20:02:06 2014, ericq, Update, Green Locking, uPDH Box Checkup 
       Reply  Wed Sep 3 00:17:22 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Wed Sep 3 20:38:40 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Fri Sep 5 04:07:44 2014, ericq, Update, Green Locking, X end uPDH Box Replaced 
                Reply  Fri Sep 5 16:13:00 2014, ericq, Update, Green Locking, new X Green PDH measurements 6x
                   Reply  Sat Sep 6 01:48:54 2014, rana, Update, Green Locking, new X Green PDH measurements 
                      Reply  Sat Sep 6 14:49:12 2014, ericq, Update, Green Locking, new X Green PDH measurements xGComparison.pdfxGspectra.pdf
                         Reply  Sun Sep 7 17:06:38 2014, rana, Update, Green Locking, new X Green PDH measurements 
                         Reply  Mon Sep 8 14:14:13 2014, Jenne, Update, Green Locking, new Y Green PDH measurements 6x
Message ID: 10464     Entry time: Sat Sep 6 14:49:12 2014     In reply to: 10463     Reply to this: 10465   10471
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: new X Green PDH measurements 

I used a modulation depth of 0.3, which, if I recall correctly, is what we aimed for on the Y-arm when we adjusted the LO signal there. However, this is probably not the case for the X arm. 

In any case, I found the bug in my RMS calculation. (I had forgotten to flip the x array in addition to the y array for the right-to-left integration, and had uneven bin spacing, so the integration bandwidths weren't correct...)

Here are the updated plots. The properly evaluated RMS is ~600Hz, which seems to mostly come in around 10k, so we may want to turn down the gain for less gain peaking in that region. 

xGComparison.pdf xGspectra.pdf

ELOG V3.1.3-